PLEASE HELP!!Consider an adiabatic constant pressure batch reactor for the following ideal gas phase reaction: 2. A—В+С The reactor initia
Answer #1

Solution - A - B + Initial Mixture c Contains 150 mola). Do 50 malus. - Ineat) Nio=0.5 , To= 300k So NA = 0.5 -8A = KCA firstFor adiabatic Reactor ä -o T-To = (-AHRT) NAC i dra Enici ENicei = NA CPA + NO Cers & Nelec & Ng CPI NA NA. (1-7A) No- NATA N